OpenStudy (anonymous):

Let cos p =.22. Which of the following is cos p/2

OpenStudy (anonymous):

half angle formula

OpenStudy (anonymous):

\[\cos(\frac{x}{2})=\pm\sqrt{\frac{1+\cos(x)}{2}}\]

OpenStudy (amistre64):

given an abc right triangle: \[cos(\frac{x}{2}) =\frac{ a}{\sqrt{2c(c-b)}}\]

OpenStudy (anonymous):

oooh what is that?

OpenStudy (amistre64):

thats just something a determined the other day :)

OpenStudy (anonymous):

where is the \(x\) in it? i mean which angle is it?

OpenStudy (amistre64):

any angle

OpenStudy (anonymous):

i cant figure it out. i suck at trig :(

OpenStudy (amistre64):

in this case; a=sqrt(1-b^2)

OpenStudy (amistre64):